OpenStudy (anonymous):

I WILL FAN AND GIVE MEDALS Find the product of (x - 7)2. x2 - 14x + 49 x2 + 14x + 49 x2 - 49 x2 + 49

OpenStudy (acxbox22):

(x-7)(x-7)=x^2-7x-7x+49....using FOIL now simplify the -7x

OpenStudy (anonymous):

doesnt the 7x cancel out right?

OpenStudy (gschibby):

No! (x-7)2 does NOT equal (x-7)(x-7)!

OpenStudy (acxbox22):

he meant  look at all the answer choices

OpenStudy (gschibby):

\[(x-7)(x-7)=(x-7)^{2}\]

OpenStudy (gschibby):

Ah, got it! ;)

OpenStudy (anonymous):

im confused